OpenStudy (anonymous):

Solve. x = –9 x = –3 x = 0 x = 4

OpenStudy (anonymous):

\[(\sqrt{6})^8x = 216^x-3\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the x is part of the exponent with 8

OpenStudy (phi):

is the -3 also part of the exponent?

OpenStudy (phi):

\[ (\sqrt{6})^{8x} = 216^{x-3} \] (put curly braces around 8x and x-3)

OpenStudy (phi):

if so, I would change the square root to a 1/2 power, and I would rewrite 216 as 6^3